During the year, a manufacturing company had the following operating results:

Beginning work-in-process inventory $ 45,000
Beginning finished goods inventory $ 190,000
Direct materials used in production $ 308,000
Direct labor $ 475,000
Manufacturing overhead incurred $ 250,000
Ending work-in-process inventory $ 67,000
Ending finished goods inventory $ 89,000


What is the cost of goods manufactured for the year?

A. $1,112,000

B. $1,134,000

C. $1,033,000

D. $1,011,000
Answer #1

Total manufacturing cost=Direct materials+Direct labor+Manufacturing overhead

=(308000+475000+250,000)=$1,033,000

Cost of goods manufactured=Total manufacturing cost+Beginning WIP-Ending WIP

=(1,033,000+45000-67000)

=$1,011,000.